Grade 11Mechanics

  1. h=2sigma*d/rgcos theta check the correctness of the equations

Mohit Gandhi
8 Years agoGrade 11

1 Answer

Eshan
8 Years ago
Dear student,

\sigmais surface tension which is force per unit length.
Hence dimensions of it are[MT^{-2}]
Hence dimensions of\dfrac{2\sigma cos\theta}{r\rho g}is\dfrac{[MT^{-2}]}{[L][ML^{-3}][LT^{-2}]}=[L]iswhich is the dimesions of h.